OpenStudy (anonymous):

Find the product of 11 terms in G.P whose 6th term is 5

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Middle term is 5 this means

ganeshie8 (ganeshie8):

wow ! yes we can use symmetry

OpenStudy (anonymous):

yesss

ganeshie8 (ganeshie8):

\[\large a_1 a_2 a_3 a_4 \cdots a_{11}\]

OpenStudy (anonymous):

A_1 * A_11 = A _2 * A _10 Like this right

ganeshie8 (ganeshie8):

that looks perfect !!

OpenStudy (anonymous):

5^11

OpenStudy (anonymous):

!!!

ganeshie8 (ganeshie8):

middle term = geometric mean

ganeshie8 (ganeshie8):

\(\large a_6 = \sqrt{a_1 \times a_{11}} = \sqrt{a_2 \times a_{10}} = \cdots \)

ganeshie8 (ganeshie8):

5^11 is correct !

OpenStudy (anonymous):

There is a formula for the product of n terms in G.P i saw it on wikipedia

OpenStudy (anonymous):

But we can't really use it here

ganeshie8 (ganeshie8):

\(\large a_1, a_1r, a_1r^2, \cdots \) product = \(\large a_1 \times a_1r \times a_1r^2 \cdots \)

ganeshie8 (ganeshie8):

if there are n terms : product = \(\large a_1^n r^{0 + 1 + 2 + \cdots n-1}\)

ganeshie8 (ganeshie8):

product = \(\large a_1^n r^{\frac{n(n-1)}{2}}\)

OpenStudy (anonymous):

yes!

ganeshie8 (ganeshie8):

nice :)

OpenStudy (anonymous):

Actually there is so much work done in this chapter especially by Euclid I felt that doing some more 15 pages i would attain a mastery at this concept , but i was stumped by reading wikipedia

OpenStudy (anonymous):

One can study a math concept till his death and still he won't learn everything

ganeshie8 (ganeshie8):

true ! cuz its so much fun that one doesn't feel like giving it up - so one pursues it till one's death xD

OpenStudy (anonymous):

hahahhah